Between 1992 and 1998 the percent of college freshman who planned to eventually get some type of medical degree can be approximated by s=-0.2369^2 + 1.425x + 6.905 where s represents the number of students and x represents the year. X=0 corresponds to 1992. In what year did this percentage reach a maximum?
Sorry, not from a text
Thanks in advance

Answer by ankor@dixie-net.com(22740)

I made the 1st term into -.2369x^2, it makes more sense that way.
:
Between 1992 and 1998 the percent of college freshman who planned to eventually get some type of medical degree can be approximated by s=-0.2369x^2 + 1.425x + 6.905 where s represents the number of students and x represents the year. X=0 corresponds to 1992. In what year did this percentage reach a maximum?
:
s = -0.2369x^2 + 1.425x + 6.905
:
Since this is a quadratic equation in the form ax^2 + bx + c, we can find the
axis of symmetry using the formula x = -b/(2a); a = -.2369, b = +1.425
:
x = -1.425/(2*-.2369)
x = -1.425/-.4738
x = +3.0
:
Substitute 3 for x in the original equation and find the vertex, that value will be the max s.
:
s = -.2369(3^2) + 1.425(3) + 6.905
:
s = -.2369(9) + 4.35 + 6.905
:
s = -2.13 + 4.35 + 6.91
:
s ~ 9 added to 1992 ~ 2001
:
Did this help?